OpenStudy (anonymous):

David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, but as he gets tired his speed decreases (at a rate of 1 mile per hour, every hour). Which equation represents the speed of the boat for x hours spent rowing? y = x + 4 y = x − 4 y = 4 − x y = 6 − x y = 6 − 2x

OpenStudy (triciaal):

i would choose y = 6-x for the first hour he did 6 miles before he got tired rate is -1/hour for x hours = -x y = mx + b y = -x + 6 answer choice rearranged